You want to install a 1 yd wide walk around a circular swimming pool. The diameter of the pool is 24 yds. What is the area of the walk?
Use 3.14 for pi.

Answer by Edwin McCravy(20055)



The green line is the radius of the swimming pool, the INNER circle.
Since its diameter is 24 yards, its radius is 12 yards.
Its area is gotten by using A = pr2.
A = 3.14×122 = 452.16 square yards.

Since the walk around the swimming pool is 1 yd. wide, the radius of
the OUTER circle has a radius of 1 yd. more than the radius of the inner 
circle, or 13 yd. Its area is also gotten by using A = pr2.
A = 3.14×132 = 530.66 square yards.

Area of walk = Area of OUTER circle MINUS Area of INNER circle:

Answer: 530.66 sq. yd. - 452.16 = 78.5 square yards. 

Edwin